# Designing lattice spin models and magnon gaps with supercurrents

### v1

@@ Abstract
Controlling magnetic interactions at the level of individual spins is relevant for a variety of quantum applications, such as qubits, memory and sensor functionality. Finding ways to exert electrical control over spin interactions, with minimal energy dissipation, is a key objective in this endeavour. We show here that spin lattices and magnon gaps can be controlled with a supercurrent. Remarkably, a spin-polarized supercurrent makes the interaction between magnetic adatoms placed on the surface of a superconductor depend not only on their relative distance, but also on their absolute position in space. This property permits electric control over the interaction not only between two individual spins, but in fact over an entire spin lattice, allowing for non-collinear ground states and a practical arena to study the properties of different spin Hamiltonians. Moreover, we show that a supercurrent controls the magnon gap in antiferromagnetic and altermagnetic insulators. These results provide an accessible way to realize electrically controlled spin switching and magnon gaps without dissipative currents.
@@
Introduction— Electrical control of magnetic interactions is relevant for a variety of quantum applications, such as sensing, computation and memory functionality. Quantum sensing with spin defects Roberts et al. (2025); Fang et al. (2024) does, for example, allow for optically measuring nuclear magnetic resonance (NMR) using the hyperfine magnetic interaction between nuclear spins and spin defects Fang et al. (2024). Quantum computing requires interactions between qubits, such as the magnetic Ruderman–Kittel–Kasuya–Yosida (RKKY) interaction connecting spin qubits Tanamoto and Ono (2021); Yang et al. (2016). Moreover, the Yu-Shiba-Rusinov states that form due to the magnetic interaction between impurity spins and electrons in a superconductor can potentially be used as a qubit Mishra et al. (2021).
@@
The RKKY interaction mediated by electrons in a superconductor provides one promising way of controlling the coupling between impurity spins Nadj-Perge et al. (2014); Pupim and Scheurer (2025); Kim et al. (2018); Heimes et al. (2015); Mohanta et al. (2018); Yazdani et al. (1997); Röntynen and Ojanen (2015). Of particular interest is typically the degree of spin non-collinearity, as this gives rise to interesting spin textures such as spin spirals Kimura (2007); Tokura and Seki (2010), skyrmions Nagaosa and Tokura (2013); Zhang et al. (2020); Marrows and Zeissler (2021) and non-collinear antiferromagnetism Rimmler et al. (2025), all with potential spintronic applications. The ground state configuration of an ensemble of spins can be non-collinear due to a Dzyaloshinskii-Moriya (DM) interaction induced by spin orbit coupling Soumyanarayanan et al. (2016) or from mixed-parity superconductivity with $s+ip_{x}$ symmetry Ouassou et al. (2025). Another quantity of interest in an ensemble of interacting spins is the magnon energy gap Chumak et al. (2015); Oba et al. (2015); Pradipto et al. (2017); Yoshii et al. (2020); Zhu et al. (2021); Yumnam et al. (2024); Mardele et al. (2024). A tunable magnon gap can, as an example, give rise to a tunable lattice thermal conductivity Vu et al. (2023). The interaction of magnons with superconducting qubits Tabuchi et al. (2015, 2016) and with flux quanta in superconductors Dobrovolskiy et al. (2019); Niedzielski et al. (2023) makes generation, classification and tuning of the magnons possible in superconducting hybrid systems.
@@
In this Letter, we show that supercurrents in superconductors enable electric control over both magnon gaps and magnetic interactions in a way that permits controllable design of a spin lattice. We use second-order perturbation theory to show analytically that a spin supercurrent carried by triplet Cooper pairs gives rise to a number of different interaction terms between magnetic adatoms placed on the superconductor. This causes the ground state configuration of the spins to be non-collinear. Importantly, we find that the ground state depends on the center-of-mass coordinate of the spins. This property allows electric control over the interaction between two spins as illustrated in Fig. 1, and in fact over an entire spin lattice, allowing for the creation of non-collinear ground states that comprise an arena to study the properties of different spin Hamiltonians. Moreover, we demonstrate that when a spin-compensated condensate of spins in the form of an antiferromagnetic or altermagnetic insulator is placed on top of the superconductor, the supercurrent controls the magnitude of the magnon gap. This can be thought of as a dissipationless magnon transistor. The result holds even for a conventional singlet BCS superconductor and occurs already at first order in perturbation theory. The electrical control of a spin lattice and the magnon gap via supercurrents predicted in this Letter reveals a synergy between superconductivity and magnetism that opens new research avenues to explore.
@@
Figure 1: Two spins $\bm{S}_{1}$ and $\bm{S}_{2}$, here represented as large arrows in the $xy$–plane, are placed on the top of a 1D superconductor with equal-spin Cooper pairs. The circles represent the lattice points, and the lattice constant is $a$. The interaction between the two spins, and hence their ground state configuration, can be tuned by a supercurrent. The positions of the spins are $r_{1},r_{2}$ and the relative coordinate is $\tau=r_{2}-r_{1}$. (a) shows the ground state configuration for two spins separated by a distance $\tau/a=1$ and $\tau/a=3$ when the supercurrent is zero. When a charge current $aQ^{\sigma}=2\pi\cdot 70/10000\approx 0.044$ is turned on in (b), the ground state changes for $\tau/a=3$. In (c), a spin current $aQ^{\sigma}=2\pi\sigma\cdot 70/10000$ flows through the superconductor. This alters the ground state compared to (a) and (b), and the spins are allowed to be non-collinear. Moreover, the spin current creates a dependence on the center-of-mass coordinate of the spins, unlike conventional spin-spin interactions that only depend on the relative coordinate. The blue arrows represent the ground state of the spins when they are moved 8 lattice sites to the right. The chemical potential is set to $\mu=-t$.
@@
Effective spin-spin interaction—For simplicity, we consider a one-dimensional (1D) superconducting chain with magnetic adatoms placed on its surface Liebhaber et al. (2022). We emphasize that our results carry over to 2D and 3D systems as well. The full Hamiltonian for the system is $H=H_{\text{SC}}+H_{c}$, where $H_{\text{SC}}$ is the Hamiltonian for the superconductor and $H_{c}$ is the coupling between the magnetic adatoms and the electrons in the superconductor. We consider a superconductor with equal-spin triplet Cooper pairs: this scenario applies both to induced triplet superconductivity via the proximity effect and to an intrinsic triplet superconductor. The former can be achieved by proximitizing an $s$-wave superconductor with one or more ferromagnets Halterman et al. (2007, 2008); Kalcheim et al. (2015) or a conductor with Rashba spin-orbit coupling Reeg and Maslov (2015), or by irradiating an $s$-wave superconductor with light Gassner et al. (2024). The starting point is the attractive Hubbard model in real space,
$\begin{split}H_{\text{SC}}=-\mu\sum_{i,\sigma}c_{i,\sigma}^{\dagger}c_{i,\sigma}-2t\sum_{\langle i,j\rangle,\sigma}c_{i,\sigma}^{\dagger}c_{j,\sigma}\\ -\frac{1}{4}U\sum_{\langle i,j\rangle}\sum_{\sigma}c_{i,\sigma}^{\dagger}c_{i,\sigma}c_{j,\sigma}^{\dagger}c_{j,\sigma}.\end{split}$ (1)
Here, $\mu$ is the chemical potential, $t$ is the hopping constant between nearest neighbors, and $U>0$ is an attractive nearest-neighbor interaction. The operators $c_{i,\sigma}^{\dagger}$ and $c_{i,\sigma}$ creates and destroys, respectively, an electron with spin $\sigma$ at lattice site $i$. $\langle i,j\rangle$ denotes a sum over nearest-neighbor pairs without double counting. We apply the mean-field approximation to the last term in $H_{\text{SC}}$. The real-space order parameter (OP) is defined as $\Delta_{i,j}^{\sigma}=-U\langle c_{j,\sigma}c_{i,\sigma}\rangle/4$, and to model a supercurrent-carrying state we write it as Takashima et al. (2017)
$\Delta_{i,j}^{\sigma}=\Delta_{\delta}^{\sigma}\text{e}^{\text{i}{Q}^{\sigma}\cdot({r}_{i}+{r}_{j})},$ (2)
@@
Here, $\Lambda$ is the $s-d$ coupling strength, $N$ is the number of lattice points, and $\bm{\sigma}$ is the Pauli vector. To derive the effective spin-spin interaction, the magnetic adatoms are treated as classical spins $\bm{S}_{i}$ at site $i$ of the superconductor.
Figure 2: The ground state configuration in the $xy$-plane of two spins $\bm{S}_{1}$ and $\bm{S}_{2}$ whose interaction is mediated by a superconductor with equal-spin Cooper pairs. Spins of the same color are separated by $\tau/a$. The position of $\bm{S}_{1}$ is given by $r_{1}/a$, and the vertical gray lines separate spin pairs with different $r_{1}/a$. The chemical potential is $\mu=-t$. (a) There are no supercurrents in the superconductor, ${Q}^{\sigma}=0$. The ground states at $\tau/a=3$ and $\tau/a=6$ are degenerate in the $xz$-plane with negligible $y$-components (not shown here). (b) A charge current $aQ^{\sigma}=2\pi\cdot 70/10000$ flows through the superconductor. (c) A spin current $aQ^{\sigma}=2\pi\sigma\cdot 70/10000$ flows through the superconductor.
We proceed to diagonalize $H_{\text{SC}}$ and perform a Schrieffer-Wolff transformation on $H$. We treat $H_{c}$ as the perturbation, and average out the fermions in the superconductor to get a pure spin Hamiltonian. The details are provided in SM . For a pure charge or spin supercurrent, there is no coupling between the magnetic adatoms and the superconductor to first order in $\Lambda.$ In this case, the effective spin Hamiltonian is of second order in $\Lambda$, and we find it to be
@@
A detailed analytical expression for the coupling constants is given in SM . They all depend on ${Q}^{\sigma}$, meaning that the spin-spin interaction can be tuned electrically via the supercurrent. The terms are interpreted as follows. $J_{i,j}$ is the exchange coupling, and usually the largest of the coupling constants. It prefers the spins to be aligned in a parallel or antiparallel configuration, depending on its sign. $K_{i,j}$ is also an exchange coupling, which tries to align the spins in a parallel or antiparallel configuration along $z$. $D_{i,j}$ is a DM interaction that prefers the spin not to be aligned along the same axis. This term is zero in the absence of a spin supercurrent. A positive $L_{i,j}$ prefers spins to be aligned along $y$ and not along $x$, and oppositely for negative $L_{i,j}$. $M_{i,j}$ tries to achieve a $\pi/2$ misalignment between the spins in the $xy$-plane, and it is zero in the absence of a spin supercurrent.
Importantly, $L_{i,j}$ and $M_{i,j}$ enable the design of a spin lattice via supercurrents. They are given by:
$L_{i,j}$ $=\frac{\Lambda^{2}}{4N^{2}}\cos\left(\Delta{Q}^{\uparrow\downarrow}\cdot({r}_{i}+{r}_{j})\right)\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}\cdot{\tau}),$
$M_{i,j}$ $=\frac{-\Lambda^{2}}{4N^{2}}\sin\left(\Delta{Q}^{\uparrow\downarrow}\cdot({r}_{i}+{r}_{j})\right)\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}\cdot{\tau}).$ (5)
@@
where $\tau=r_{j}-r_{i}$ while the coefficients $A_{q,\downarrow,\uparrow}^{(2)}$ are large expressions that depend on energy eigenvalues and eigenstates of the superconducting Hamiltonian as well as temperature. Their dependence on the absolute position $r_{i}+r_{j}$ in the presence of a spin current $\Delta Q^{\uparrow\downarrow}={Q}^{\uparrow}-{Q}^{\downarrow}$ stands in stark contrast to the conventional relative position dependence $\tau$ that usually appears for RKKY interactions. This is a remarkable feature that arises solely in the presence of a spin supercurrent. It is this dependence that enables the electrical design of a spin lattice, as we will proceed to show. The dependence on absolute position can be understood physically as follows. A Cooper pair of spin $\uparrow$ electrons with one electron located at $r_{i}$ and the other electron at $r_{j}$ has the phase $Q^{\uparrow}\cdot(r_{i}+r_{j})$, as seen from Eq. (2). This phase is arbitrary since the physics cannot change under a U(1) gauge transformation of the order parameter. However, phase differences have physical consequences, and in the presence of a spin supercurrent, such a phase difference exists between a Cooper pair with spin $\uparrow$ and a Cooper pair with spin $\downarrow$. This phase difference is $\Delta Q^{\uparrow\downarrow}\cdot(r_{i}+r_{j})$, which is precisely the quantity that the spin interaction coefficients depend on.
[…]
To illustrate how the ground state spin configuration is altered by a supercurrent, we consider two magnetic adatoms $\bm{S}_{1}$ and $\bm{S}_{2}$ placed on top of a superconductor with equal-spin Cooper pairs. The position of $\bm{S}_{1}$ is $r_{1}$, and the position of $\bm{S}_{2}$ is $r_{1}+\tau$. Figure 2(a) shows the ground state of these two spins in the absence of a supercurrent for various $r_{1}$ and $\tau$ values. The ground state configuration can be ferromagnetic or antiferromagnetic, depending on the distance between the spins. Upon applying a charge supercurrent, shown in Fig. 2(b), the spin alignment axis changes for instance from the $xz$-plane to the $y$-axis for $\tau/a=3$. Since there is no spin current, the ground state configuration depends on the relative position of the spins $\tau$, but not on the absolute position of the spins. This changes when the supercurrent is spin-polarized. In Fig. 2(c), a spin supercurrent flows through the superconductor. The onset of non-zero $D_{i,j}$ and $M_{i,j}$ causes the spins to be non-collinear, and the angle difference from a collinear state is up to 65 degrees. Moreover, the ground state configuration now depends on the absolute position of the spins in the lattice. This enables the possibility to design a lattice spin model that breaks translational invariance. We have verified that these properties are robust upon varying the chemical potential SM .
Magnon gap
@@
We now replace the two impurity spins considered up to now with an entire spin condensate that is proximity-coupled to a superconductor to demonstrate a new effect: electric control of a magnon gap in a spin-compensated magnet via a supercurrent. Consider a bilayer consisting of an antiferromagnetic insulator and a superconductor with equal-spin Cooper pairs relative to the Néel order, as illustrated in Fig. 3(a). The first order contribution to the magnon gap in the antiferromagnetic insulator is found by calculating $\langle H_{c}\rangle_{\gamma}$ where $\gamma$ are the fermionic operators that diagonalize $H_{\text{SC}}$ , applying the Holstein-Primakoff transformation to the spin operators, and diagonalizing the full Hamiltonian for the antiferromagnetic insulator. We find that $\langle H_{c}\rangle_{\gamma}$ is equivalent to an applied magnetic field, and the analytical supercurrent-induced contribution to the magnon gap is
$\mathcal{D}$ $=-\frac{\Lambda}{2N}\sum_{k,\sigma}\sigma\big{(}|u_{k,\sigma}|^{2}n(E_{k+Q^{\sigma},\sigma})$
$+|v_{k,\sigma}|^{2}[1-n(E_{-k+Q^{\sigma},\sigma})]\big{)},$ (6)
@@
where $n$ is the Fermi-Dirac distribution. The calculation details as well as definitions of the eigenenergy $E_{k,\sigma}$ and the coherence factors $u_{k,\sigma}$ and $v_{k,\sigma}$ are provided in SM . A pure charge supercurrent $(Q_{\uparrow}=Q_{\downarrow})$ or a pure spin supercurrent $(Q_{\uparrow}=-Q_{\downarrow})$ gives $\mathcal{D}=0$ , while any combination of a charge and spin supercurrent $(Q_{\uparrow}\neq\pm Q_{\downarrow})$ gives a tunable $\mathcal{D}\neq 0$. This is shown in Fig. 3(b). In other words, a spin-polarized supercurrent tunes the magnitude of the magnon gap. When $\mathcal{D}>0$, the magnon gap is enhanced for one magnon species and suppressed for the other magnon species.
@@
Interestingly, there may exist an even simpler way to control magnon gaps in antiferromagnetic insulators experimentally. Consider a bilayer consisting of a thin-film conventional $s$-wave superconductor with a spin-splitting field and an antiferromagnetic insulator on top, as illustrated in Fig. 3(c). The spin-splitting field in the superconductor can be achieved either via an applied in-plane external field or by proximity to a ferromagnet underneath the superconductor. The supercurrent-induced contribution to the magnon gap is analytically computed to be SM :

### v2

@@ Abstract
Electric control over magnetic interactions at the level of individual spins is relevant for a variety of quantum applications, such as qubits, memory and sensor functionality. We show here that spin lattices and magnon gaps can be controlled with a supercurrent. Remarkably, a spin-polarized supercurrent makes the interaction between magnetic adatoms placed on the surface of a superconductor depend not only on their relative distance, but also on their absolute position in space. This property permits electric control over the interaction not only between two individual spins, but over an entire spin lattice, allowing for tunable non-collinear ground states and a practical arena to study the properties of different spin Hamiltonians. Moreover, we show that a supercurrent controls the magnon gap in antiferromagnetic and altermagnetic insulators. These results provide an accessible way to realize electrically controlled spin switching and magnon gaps without dissipative currents.
@@
Introduction— Electrical control of magnetic interactions is relevant for a variety of quantum applications, such as sensing, computation and memory functionality. Quantum sensing with spin defects Roberts et al. (2025); Fang et al. (2024) allows for optically measuring nuclear magnetic resonance (NMR) Fang et al. (2024). Quantum computing requires interactions between qubits, such as the magnetic Ruderman–Kittel–Kasuya–Yosida (RKKY) interaction connecting spin qubits Tanamoto and Ono (2021); Yang et al. (2016); Mishra et al. (2021). Finally, memory functionality based on electric writing of the spin-state improves the efficiency of such devices Zhang et al. (2024).
@@
The RKKY interaction mediated by electrons in a superconductor provides one promising way of controlling the coupling between spins at an individual level Nadj-Perge et al. (2014); Pupim and Scheurer (2025); Kim et al. (2018); Heimes et al. (2015); Mohanta et al. (2018); Yazdani et al. (1997); Röntynen and Ojanen (2015). Of particular interest is typically the degree of spin non-collinearity, as this gives rise to interesting spin textures such as spin spirals Kimura (2007); Tokura and Seki (2010), skyrmions Nagaosa and Tokura (2013); Zhang et al. (2020); Marrows and Zeissler (2021) and non-collinear antiferromagnetism Rimmler et al. (2025), all with potential spintronic applications. The ground state configuration of an ensemble of spins can be non-collinear due to a Dzyaloshinskii-Moriya (DM) interaction induced by spin orbit coupling Soumyanarayanan et al. (2016) or from mixed-parity superconductivity with $s+ip_{x}$ symmetry Ouassou et al. (2025). Another quantity of interest in an ensemble of interacting spins is the magnon energy gap Chumak et al. (2015); Oba et al. (2015); Pradipto et al. (2017); Yoshii et al. (2020); Zhu et al. (2021); Yumnam et al. (2024); Mardele et al. (2024). A tunable magnon gap can, as an example, give rise to a tunable lattice thermal conductivity Vu et al. (2023). The interaction of magnons with superconducting qubits Tabuchi et al. (2015, 2016) and flux quanta in superconductors Dobrovolskiy et al. (2019); Niedzielski et al. (2023) makes the generation and tuning of magnons possible in superconducting hybrid systems.
@@
In this Letter, we show that supercurrents in superconductors enable electric control over both magnon gaps and magnetic interactions in a way that permits controllable design of a spin lattice. We use perturbation theory to show analytically that a spin supercurrent carried by triplet Cooper pairs gives rise to a number of different interaction terms between magnetic adatoms placed on the superconductor Liebhaber et al. (2022). This causes the ground state configuration of the spins to be non-collinear. Importantly, we find that the ground state depends on the center-of-mass coordinate of the spins, unlike the usual RKKY interaction. This property allows control over the interaction between two spins as illustrated in Fig. 1, and in fact over an entire spin lattice, providing an arena to study the properties of different spin Hamiltonians. Moreover, we demonstrate that when an antiferromagnetic or altermagnetic insulator is placed on top of the superconductor, the supercurrent controls the magnitude of the magnon gap. This can be thought of as a dissipationless magnon transistor and is achieved even with a conventional singlet BCS superconductor. The magnon gap correction is a first-order effect in the coupling between spins and quasiparticles in the superconductor, while the spin-spin interactions enabling spin lattice design appear at second order. The electrical control of a spin lattice and the magnon gap via supercurrents predicted in this Letter reveals a synergy between superconductivity and magnetism that opens new research avenues to explore.
@@
Figure 1: Two spins $\bm{S}_{1}$ and $\bm{S}_{2}$, here represented as large arrows in the xyxy–plane, are placed on the top of a 1D superconductor with equal-spin Cooper pairs. The circles represent the lattice points, and the lattice constant is $a$. The interaction between the two spins, and hence their ground state configuration, can be tuned by a supercurrent. The positions of the spins are $r_{1},r_{2}$ and the relative coordinate is $\tau=r_{2}-r_{1}$. (a) shows the ground state configuration for two spins separated by a distance $\tau/a=1$ and $\tau/a=3$ when the supercurrent is zero. When a charge current $Q^{\sigma}/Q_{c}\approx 0.87$ is turned on in (b), the ground state changes for $\tau/a=3$. In (c), a spin current $Q^{\sigma}/Q_{c}\approx 0.87\sigma$ flows through the superconductor. This alters the ground state compared to (a) and (b), and the spins are allowed to be non-collinear. Moreover, the spin current creates a dependence on the center-of-mass coordinate of the spins, unlike conventional spin-spin interactions that only depend on the relative coordinate. The blue arrows represent the ground state of the spins when they are moved 8 lattice sites to the right. The chemical potential is set to $\mu=-t$. The supercurrent is normalized on the critical supercurrent $aQ_{c}=0.0505$, which is found by solving the gap equation.
@@
Effective spin-spin interaction—For simplicity, we consider an effective one-dimensional (1D) superconducting chain with magnetic adatoms placed on its surface, as this will permit us to numerically simulate large systems and since experiments often can be modeled well by effective 1D Hamiltonians Liebhaber et al. (2022); le Sueur et al. (2008). Our results carry over straightforwardly to 2D and 3D systems SM . The full Hamiltonian for the system is $H=H_{\text{SC}}+H_{c}$, where $H_{\text{SC}}$ is the Hamiltonian for the superconductor and $H_{c}$ is the coupling between the magnetic adatoms and the electrons in the superconductor. We consider a superconductor with equal-spin triplet Cooper pairs: this scenario applies both to induced triplet superconductivity via the proximity effect and to an intrinsic triplet superconductor. The former can be achieved by proximitizing an $s$-wave superconductor with one or more ferromagnets Halterman et al. (2007, 2008); Kalcheim et al. (2015); Grein et al. (2013); Eschrig and Löfwander (2008); Khaire et al. (2010); Robinson et al. (2010) or a conductor with Rashba spin-orbit coupling Reeg and Maslov (2015), curving the superconductor Ying et al. (2017), or by irradiating it with light Gassner et al. (2024). The starting point is the attractive extended Hubbard model in real space,
$H_{\text{SC}}=-\mu\sum_{i,\sigma}n_{i,\sigma}-\sum_{\langle i,j\rangle,\sigma}\Big(2tc_{i,\sigma}^{\dagger}c_{j,\sigma}+\frac{U}{4}n_{i,\sigma}n_{j,\sigma}\Big)$ (1)
@@
Here, $\Lambda$ is the sds-d coupling strength, and $\bm{\sigma}$ is the Pauli vector. Physically, the exchange $\Lambda$ can be modified by changing the overlap between the wavefunctions of the quasiparticles and the magnetic adatoms. To derive the effective spin-spin interaction, the magnetic adatoms are treated as classical spins $\bm{S}_{i}$ at site $i$ of the superconductor.
Figure 2: (a-d) The ground state configuration in the $xy$-plane of two spins $\bm{S}_{1}$ and $\bm{S}_{2}$ whose interaction is mediated by a superconductor with equal-spin Cooper pairs. Spins of the same color are separated by $\tau/a$. (a) There are no supercurrents in the superconductor, ${Q}^{\sigma}=0$. The ground states at $\tau/a=3$ and $\tau/a=6$ are degenerate in the $xz$-plane with negligible $y$-components (not shown here). (b) A charge current $Q^{\sigma}/Q_{c}\approx 0.87$ flows through the superconductor. In both (a) and (b), there is no center-of-mass-dependence $(r_{1})$. (c) The position of $\bm{S}_{1}$ is given by $r_{1}/a$, and the vertical gray lines separate spin pairs with different $r_{1}/a$. A spin supercurrent $Q^{\sigma}/Q_{c}\approx 0.87\sigma$ flows through the superconductor and makes the RKKY interaction depend on the center-of-mass coordinate of the two spins. (d) Spin supercurrent dependence of the magnetic ground state for $r_{1}/a=8$. (e) Spin supercurrent dependence for a trimer with spin positions $r_{1}/a=4$, $r_{2}/a=6$ and $r_{3}/a=8.$ The chemical potential is $\mu=-t$.
We proceed to first diagonalize the mean-field Hamiltonian $H_{\text{SC}}$ and then perform a Schrieffer-Wolff transformation on $H$. We treat $H_{c}$ as the perturbation, and average out the fermions in the superconductor to get a pure spin Hamiltonian SM . For a pure charge or spin supercurrent, there is no coupling between the magnetic adatoms and the superconductor to first order in $\Lambda.$ In this case, the effective spin Hamiltonian is of second order in $\Lambda$, and we find it to be
@@
A detailed analytical expression for the coupling constants is given in SM . They all depend on ${Q}^{\sigma}$, meaning that the spin-spin interaction can be tuned electrically via the supercurrent. Since $\mathcal{H}_{\text{eff}}$ is obtained via perturbation theory, the expression is valid when the energy scale of the predicted effects is small compared to the terms in the original unperturbed Hamiltonian. We have confirmed numerically that the coupling constants are small compared to $\Delta_{0},t$, and $\mu$. In $\mathcal{H}_{\text{eff}}$, $J_{i,j}$ and $K_{i,j}$ are exchange couplings, and the DM term $D_{i,j}$ is nonzero only in the presence of a spin supercurrent. Importantly, $L_{i,j}$ and $M_{i,j}$ enable the design of a spin lattice via supercurrents. They are given by
$L_{i,j}=\Upsilon\cos\left(\Delta{Q}^{\uparrow\downarrow}R\right),\;M_{i,j}=-\Upsilon\sin\left(\Delta{Q}^{\uparrow\downarrow}R\right),$ (4)
@@
where $\Upsilon\equiv\frac{\Lambda^{2}}{4N^{2}}\sum_{{q}}A^{(2)}_{{q},\downarrow,\uparrow}\cos({q}{\tau})$, $\Delta{Q}^{\uparrow\downarrow}=Q^{\uparrow}-Q^{\downarrow}$, $R=r_{i}+r_{j}$, $\tau=r_{j}-r_{i}$, and the coefficients $A_{q,\downarrow,\uparrow}^{(2)}$ are large expressions that depend on energy eigenvalues and eigenstates of the superconducting Hamiltonian. The dependence of $L_{i,j}$ and $M_{i,j}$ on the absolute position $r_{i}+r_{j}$ stands in stark contrast to the conventional relative position dependence $\tau$ that usually appears for RKKY interactions. This is a remarkable feature that arises solely in the presence of a spin supercurrent. It is this dependence that enables the electrical design of a spin lattice, as we will show in the next section, and it can be understood physically as follows. A Cooper pair of spin $\uparrow$ electrons with one electron located at $r_{i}$ and the other electron at $r_{j}$ has the phase $Q^{\uparrow}R$. This phase is arbitrary since the physics cannot change under a U(1) gauge transformation of the order parameter. However, phase differences have physical consequences, and in the presence of a spin supercurrent, such a phase difference exists between a Cooper pair with spin $\uparrow$ and a Cooper pair with spin $\downarrow$. This phase difference is $\Delta Q^{\uparrow\downarrow}R$, which is precisely the quantity that the spin interaction coefficients depend on. The origin $r_{i}=0$ is defined by the relative phase $\Delta Q^{\uparrow\downarrow}$ of the superconducting condensates SM .
We expect that the functional form of $\mathcal{H}_{\text{eff}}$ would remain unchanged in the strong coupling limit, for quantum spin systems, and if taking into account the local suppression of the magnitude of the OP, although the magnitude of the coupling constants would change. We also expect the this in the presence of spin-dependent terms in the normal state Hamiltionan, such as spin-orbit coupling or spin splitting, and if an intrinsic interaction between the magnetic adatoms is present in the first place. However, any such intrinsic exchange coupling would fall off extremely rapidly upon separating the impurity spins due to the lack of wavefunction overlap, leaving only the interaction mediated by the superconductor. In the limit of large $S=|\bm{S}_{i}|$ and small coupling $\Lambda$ that we consider, the impurity spins may indeed be treated as classical and any Yu-Shiba-Rusinov (YSR) states would reside right at the gap edge and provide a negligible spin screening Balatsky et al. (2006) and negligible contribution to the spin-spin interaction Deb et al. (2021). Screening currents induced from the impurity spins can be neglected for a superconducting film of thickness $d\ll\lambda_{L}$ where $\lambda_{L}$ is the penetration depth Meservey and Tedrow (1994), as is standard in the treatment of magnetic adatoms on superconducting surfaces Lo Conte et al. (2024).
@@
Electrical tuning of the ground state —The ground state configuration is found by minimizing $\mathcal{H}_{\text{eff}}$ numerically. Note that $\mathcal{H}_{\text{eff}}$ is invariant under $\bm{S}_{i}\rightarrow-\bm{S}_{i}$, meaning that the ground state is always twofold degenerate. To illustrate how the ground state spin configuration is altered by a supercurrent, we consider two magnetic adatoms $\bm{S}_{1}$ and $\bm{S}_{2}$ placed on top of a superconductor with equal-spin Cooper pairs. The position of $\bm{S}_{1}$ is $r_{1}$, and the position of $\bm{S}_{2}$ is $r_{1}+\tau$. Figure 2(a) shows the ground state of these two spins in the absence of a supercurrent for various $\tau$ values. The ground state configuration can be ferromagnetic or antiferromagnetic, depending on the distance between the spins. Upon applying a charge supercurrent, shown in Fig. 2(b), the spin alignment axis changes for instance from the $xz$-plane to the $y$-axis for $\tau/a=3$. Since there is no spin current, the ground state configuration depends on the relative position of the spins $\tau$, but not on the absolute position of the spins. This changes when the supercurrent is spin-polarized. In Fig. 2(c), a spin supercurrent flows through the superconductor, and we observe that the ground state configuration now depends on the absolute positions $r_{1}$ and $r_{1}+\tau$ of the spins in the lattice. This enables the possibility to design a lattice spin model that breaks translational invariance. The onset of non-zero $D_{i,j}$ and $M_{i,j}$ causes the spins to be non-collinear, and the angle difference from a collinear state is up to 65 degrees. Figure 2(d) shows how the ground state can be tuned by a spin supercurrent. We have verified that all these properties are robust upon varying the chemical potential SM . Moreover, we have verified that these properties also hold for a spin trimer consisting of three classical spins SM . In Fig. 2(e), we show how the trimer ground state varies with the spin supercurrent magnitude.
We note in passing that while the ground state configuration of the magnetic adatoms could be calculated non-perturbatively by diagonalizing numerically the Bogoliubov-de-Gennes Hamiltonian Christensen et al. (2016); Pientka et al. (2013), our perturbative approach allows us to extract an analytical effective spin-spin Hamiltonian in Eq. (S13). This determines the physical origin of the center-of-mass dependence as well as the precise analytical conditions for when the coupling constants are nonzero. Going beyond our perturbative approach to a regime where YSR states can reside deep within the gap could be an interesting future venue to explore emergent topological superconductivity by manipulating the state of a magnetic adatom lattice with supercurrents Pientka et al. (2013); Awoga et al. (2024); Li et al. (2016).
@@
Magnon gap—We now replace the two impurity spins considered up to now with an entire spin condensate that is proximity-coupled to a superconductor to demonstrate a new effect: electric control of a magnon gap in a spin-compensated magnet via a supercurrent. Consider a bilayer consisting of an antiferromagnetic insulator and a superconductor with equal-spin Cooper pairs relative to the Néel order, as illustrated in Fig. 3(a). We underline that our results apply to an altermagnetic insulator as well by generalizing the model to 2D SM . The first order contribution to the magnon gap in the antiferromagnetic insulator is found by calculating $\langle H_{c}\rangle_{\gamma}$ where $\gamma$ are the fermionic operators that diagonalize $H_{\text{SC}}$ , applying the Holstein-Primakoff transformation to the spin operators, and diagonalizing the full Hamiltonian for the antiferromagnetic insulator. We find that $\langle H_{c}\rangle_{\gamma}$ is equivalent to an applied magnetic field, and the analytical supercurrent-induced contribution to the magnon gap is
$\mathcal{D}=-\frac{\Lambda}{2N}\sum_{k,\sigma}\sigma\big(|u_{k,\sigma}|^{2}n^{+}_{k,\sigma}+|v_{k,\sigma}|^{2}n^{-}_{k,\sigma}\big),$ (5)
@@
where $n^{\pm}_{k,\sigma}=n(\pm E_{\pm k+Q^{\sigma},\sigma})$ is the Fermi-Dirac distribution. The calculation details as well as definitions of the eigenenergy $E_{k,\sigma}$ and the coherence factors $u_{k,\sigma}$ and $v_{k,\sigma}$ are provided in SM . A pure charge supercurrent $(Q_{\uparrow}=Q_{\downarrow})$ or a pure spin supercurrent $(Q_{\uparrow}=-Q_{\downarrow})$ gives $\mathcal{D}=0$ , while any combination of a charge and spin supercurrent $(Q_{\uparrow}\neq\pm Q_{\downarrow})$ gives a tunable $\mathcal{D}\neq 0$. This is shown in Fig. 3(b). In other words, a spin-polarized supercurrent tunes the magnitude of the magnon gap. The magnon gap is enhanced for one magnon species and suppressed for the other magnon species.
@@
Interestingly, there may exist an even simpler way to control magnon gaps in antiferromagnetic insulators experimentally. Consider a bilayer consisting of a thin-film conventional $s$-wave superconductor with a spin-splitting field Bergeret et al. (2018) and an antiferromagnetic insulator on top, as illustrated in Fig. 3(c). The spin-splitting field in the superconductor can be achieved via an applied in-plane external field or by proximity to a ferromagnet when the thickness of the film is smaller than the London penetration depth. The induced spin splitting should be smaller than the critical value $h_{c}=\Delta_{0}/\sqrt{2}$ where $\Delta_{0}$ is the OP magnitude at $T=0$ Clogston (1962); Chandrasekhar (1962). The supercurrent-induced change in the magnon gap is given by SM :